### v1

# RationalVLA: A Rational Vision-Language-Action Model with Dual System

Abstract
A fundamental requirement for real-world robotic deployment is the ability to understand and respond to natural language instructions. Existing language-conditioned manipulation tasks typically assume that instructions are perfectly aligned with the environment. This assumption limits robustness and generalization in realistic scenarios where instructions may be ambiguous, irrelevant, or infeasible. To address this problem, we introduce RAtional MAnipulation (RAMA), a new benchmark that challenges models with both unseen executable instructions and defective ones that should be rejected. In RAMA, we construct a dataset with over 14,000 samples, including diverse defective instructions spanning six dimensions: visual, physical, semantic, motion, safety, and out-of-context. We further propose the Rational Vision-Language-Action model (RationalVLA). It is a dual system for robotic arms that integrates the high-level vision-language model with the low-level manipulation policy by introducing learnable latent space embeddings. This design enables RationalVLA to reason over instructions, reject infeasible commands, and execute manipulation effectively. Experiments demonstrate that RationalVLA outperforms state-of-the-art baselines on RAMA by a 14.5% higher success rate and 0.94 average task length, while maintaining competitive performance on standard manipulation tasks. Real-world trials further validate its effectiveness and robustness in practical applications. Our project page is https://irpn-eai.github.io/rationalvla.
Index Terms:
Robot Manipulation, Vision-Language-Action Model, Dual System
I Introduction
“Half of the troubles of this life can be traced to saying yes too quickly and not saying no soon enough.”
— Josh Billings
Embodied intelligence is considered as the ultimate form of artificial intelligence (AI) [1]. A necessary condition for the successful deployment of embodied intelligence in the real world is their ability to understand natural language and respond appropriately, either by providing correct answers or by executing the corresponding actions. This demand has sparked research on language-conditioned manipulation tasks, which require robots to follow natural language instructions to complete specific manipulation actions. Recent works have identified and investigated several key challenges in this field, particularly in handling complex and previously unseen instructions [2, 3, 4]. They primarily leverage the language generalization capabilities of pretrained vision-language models (VLMs) [5, 6, 7, 8, 9, 10], fine-tuning them on robotic datasets to obtain vision-language-action (VLA) models [11, 2, 12, 13, 14, 15], which enables the handling of various instructions to some extent.
Figure 1: Overview of two instructions. This figure illustrates two failure modes of traditional VLA models under unconventional instructions. (1) Complex and unseen instructions. Traditional manipulation policies and VLAs [16, 3] own limited language generalization abilities, thus executing wrong actions. (2) Defective instructions. Traditional VLA models have largely lost their language reasoning ability and tend to overfit to incorrect actions. In contrast, our RationalVLA successfully rejects such instructions by first perceiving the environment and then correctly interpreting the language.
However, another common and meaningful scenario in real-world applications is often overlooked: defective instructions. Defective instructions, such as instruction-environment mismatches, out-of-scope instructions, and malicious or irrelevant commands, are flawed and infeasible for the current scene. Fig. 1 shows that current policies and VLAs [16, 3] struggle with these defective instructions. If a robot attempts to execute these defective instructions, it may lead to unintended outcomes such as disrupting environmental order and causing the following task failures. This highlights the urgent need for a benchmark that reflects more generalized and realistic scenarios, where both seen, unseen, and defective instructions coexist.
To better assess models’ language understanding and generalization abilities, we propose Rational Manipulation benchmark (RAMA). RAMA benchmark comprises both executable but unseen instructions and defective instructions that do not match any executable task in the current scene. This improvement is more consistent with real-world conditions. Notably, even a small number of defective instructions can impact the completion of a long-horizon task. To support the training and evaluation on the RAMA benchmark, we established a new dataset with 14k defective samples, including 1k only for testing. The evaluation protocol complements CALVIN [17] with a prefix defective task in each rollout, and all other tasks are annotated with complex and unseen instructions [3] to further challenge language understanding.
To handle RAMA benchmark, models must possess two key abilities: understanding and reasoning on various instructions, and executing manipulation effectively. We survey two different common architectures to achieve this and analyze their deficiencies: (1) Since VLA models inherit the language understanding capabilities of pretrained VLMs, a fine-tuned VLA can, to some extent, perform both instruction comprehension and action execution. However, fine-tuning on robot action datasets often results in catastrophic forgetting [18] of language abilities, and the resulting manipulation performance still tends to be inferior to that of task-specific policies. (2) Another straightforward way to leverage VLM’s language abilities and task-specific policies’ manipulation abilities is to combine them together. This decoupled architecture can identify defective instructions. However, its language capabilities are limited in VLM, and VLM can not transmit complex language comprehension and reasoning to policies. As a result, this architecture still performs poorly when faced with complex and previously unseen instructions.
Following the above analysis, we propose the Rational Vision-Language-Action model (RationalVLA). As shown in Fig. 5, RationalVLA is a dual-system vision-language-action model, composed of a high-level MLLM and a low-level robot policy in an end-to-end training manner. Our method incorporates learnable latent space embeddings as the interface between high level and low level. RationalVLA achieves this by learning two types of tokens at the interface layer: (1) <ACT> token to instruct the low-level policy to output actions. This allows RationalVLA to inherit the perceiving and reasoning capabilities and also preserve the manipulation skills. (2) <REJ> token. Inspired by previous works on dynamically processing neural network inputs [19, 20], <REJ> token allows RationalVLA to reject the defective instructions according to current observation.
Benefiting from the novel designs, RationalVLA takes a big step forward in addressing RAMA challenges. Experiments show that RationalVLA performs better than other baselines by 14.5% on the success rate of the last tasks and 0.94 in average length. In addition, it also maintains competitive performance on regular manipulation tasks. Finally, we validate the effectiveness of our RationalVLA in the real world.
Our contributions mainly lie in three aspects:
  • We introduce a new benchmark, the Rational Manipulation (RAMA) with defective instructions in 6 dimensions, making manipulation tasks more flexible and practical for real-world deployment. We establish a dataset, which consists of more than 14k data samples with defective instructions for training and evaluation.
  • We propose Rational Vision-Language-Action model (RationalVLA). It is a dual system for the robotic arm to perceive the environment, handle unseen and defective instructions effectively, and demonstrate performant manipulation capabilities. Our extensive experiments show that it outperforms all baselines on both RAMA benchmark and classic manipulation tasks, showing its strong language generalization capabilities.
II Related Work
While previous research[21, 22, 23, 24, 25] has made notable progress in enabling robots to perform precise and efficient manipulation, our work specifically focuses on the challenges of language-conditioned manipulation. To enhance the robustness of instruction understanding in such robots, prior research relevant to our proposed RationalVLA model can be organized into three key areas: (1) language-conditioned manipulation, which focuses on learning action policies from natural language. (2) dual-system VLA models, which decouple high-level reasoning from low-level control. and (3) efforts addressing defective instructions in human-robot collaboration of RAMA benchmark. Below, we briefly review representative work in each area.
II-A Language-conditioned Manipulation
Early works [26, 27, 11] use text encoders to map language instructions to latent features, which are then conditioned by robot policies to predict actions. Some approaches [28, 29, 16, 30] employ diffusion models as action policies. RT-2 [2] and subsequent works [31, 32] integrate large language models to enhance language generalization. Inner Monologue [33] enriches instruction processing and planning in robotic control. RoboFlamingo [3] fine-tunes vision-language models for action prediction, while GR-1 [34] learns from action videos and Vision Language Planning [35] generate futural action videos from language to predict action. RoboMamba [36] utilizes state space models for efficient instruction-based reasoning. VLAS [37] extends the text language to speech. Our model demonstrates superior reasoning capabilities over these methods, enabled by its effective understanding of the intricate relationship between language and the environment. Other work [38] is a dual-arm robot framework combining dynamic manipulation primitives efficiently unfolds garments using self-supervised learning, achieving high coverage without expert data.
II-B Dual-system VLA Models
LCB [39] introduces latent codes as an alternative to natural language between LLM planners and robot controllers, enabling better expression of complex tasks and end-to-end fine-tuning. HiRT [40] introduces a hierarchical transformer that reduces reliance on heavy VLMs by combining low-frequency semantic processing with high-frequency vision-based control, enabling real-time performance on dynamic tasks. RoboDual [41] combines a generalist VLA policy for broad understanding with a lightweight specialist for efficient, high-frequency control, achieving strong generalization and real-world performance with minimal data and compute. DP-VLA [42] builds on OpenVLA [43] by using it as a frozen high-level module for reasoning, while introducing a lightweight controller for real-time execution, achieving efficient and scalable robotic manipulation. GR00T N1 [44] is an open foundation VLA model for humanoid robots, featuring a dual-system architecture that integrates high-level vision-language reasoning with real-time motor control. OpenHelix [45] uses LLaVA [5] as the MLLM and leverages prompt tuning with an auxiliary task to align its outputs with a pretrained 3D diffusion policy, enabling efficient multimodal reasoning and control without finetuning. Meanwhile, other work [46] proposes a hierarchical framework: a high-level module that corrects end-effector trajectories using dynamic movement primitives, and a low-level safety filter that filters out unsafe actions and generates safe ones via a quadratic programming method. However, these approaches lack robustness to defective instructions, which is insufficient for tackling challenges in RAMA benchmark.
II-C Defective Instructions in Human-robot Interaction
Ob-VLN [47] first considers the defective instructions in vision-language navigation. The discrepancies in actual scenes and given instructions can cause major failures. This work constructs a virtual graph to help the robot manage it. BadNaver [48] further identifies defective instructions as a type of jailbreak attack, which leads the model to execute vulnerable actions. Experiments on various robot agents imply that these instructions affect the safety. These works emphasize the importance of dealing with defective instructions. However, these works limit the question in navigation tasks, overlooking its importance in manipulation tasks.
III Rational Manipulation Benchmark
In this section, we introduce our Rational Manipulation (RAMA) benchmark.
III-A Task Setting
The RAMA benchmark is designed to evaluate the generalization and robustness of language understanding abilities. Specifically, we incorporate defective instructions to evaluate scene perception and language reasoning abilities, and complex, unseen instructions to assess the generalization capability of language understanding. Here, defective instructions refer to those that do not match any executable task in the scene. Specifically, we consider the defective instructions across six dimensions - visual, physical, semantic, motion, safety, and out-of-context. We summarize our taxonomy in Fig. 2:
  • Visual. The visual non-executable task refers to the target with non-existent visual characteristics (color, texture, and surface) in the scenario. These visual perturbations challenge the extraction and understanding of the environment.
  • Physical. The physical non-executable task refers to the target with non-existent shape and size in the scenario. The understanding of these spatial positions sets requirements for physical prior knowledge of the robot assistant.
  • Semantic. The semantic non-executable task refers to the non-existent object in the scenario. This requires the robot assistant to understand the semantics of the target object and determine whether it is present in the scene. Therefore, it challenges the rational reasoning capability of the language model.
  • Motion. The motion non-executable task refers to actions that can not be executed due to the limitations of the structure and kinematic characteristics of the mechanical arm. This requires the robot assistant to assess the feasibility of executing the action based on its capabilities.
  • Safety. Safety considerations are critically important in the field of robotics, particularly for fully autonomous robot assistants endowed with independent decision-making capabilities. More research focuses on safe interaction between operators and collaborative robots [49]. Safety problems may encompass explicit attacks and potential risks posed by robots to humans, the environment, and themselves. Safety concerns are key determinants for developing a real-world robot assistant.
  • Out-of-context. In addition to the above dimensions, some completely unrelated, illogical, and nonsensical commands should also be included as non-executable instructions to ensure the robustness robot assistant against noise.
On the other hand, we use the instructions generated by GPT-4 [50] in RoboFlamingo [3] to serve as the complex, unseen instructions. They are paraphrased versions of training instructions, ensuring correspondence to actions in the training dataset.
III-B Simulation Environment
To establish a unified and convenient benchmark, we construct a dataset and evaluation protocol based on CALVIN, which is well-regarded in language-conditioned manipulation research. 1) It ensures fair comparisons between different models and reproducibility of experiments. 2) It supports long-horizon tasks, which is crucial for evaluating the influence of defective instructions.
CALVIN is built on top of the PyBullet [51] simulator. It involves a desk with some objects and a Franka Panda Robot arm that manipulates the scene. It consists of 34 tasks and 4 different environments (A, B, C and D).
Figure 2: In RAMA benchmark, we assess models across six dimensions: visual, physical, semantic, motion, safety, and out-of-context.
III-C Dataset
We construct a large-scale, diverse dataset based on CALVIN and consider the expression of defective instructions. RAMA benchmark contains 14,412 language instructions in total, and quantitative statistics can be found in Table I. Among them, 159 instructions are held out exclusively as the test set, separated from the 14,253 training instructions, in order to evaluate the model’s language generalization capability. Inspired by [26], we utilize a hindsight approach to collect this dataset, as shown in Fig. 3.
TABLE I: The statistical details of RAMA benchmark across 6 dimensions.
Data splits Vis. Phy. Sem. Mot. Saf. Out. Mix. Total Train 1452 1302 1205 756 1085 1140 7313 14,253 Test 20 18 40 26 25 30 159 Total 1472 1320 1245 782 1110 1170 7313 14,412
Figure 3: Hindsight data collection after-the-fact. We relabeled the observations and actions in CALVIN with defective instructions after the fact. These defective instructions are generated by our customized instruction generator, either programmatically through a modular system or directly via prompts to GPT-4o.
For visual, physical, semantic, and motion defective instructions, we first identify existing targets in the CALVIN environment and then programmatically replace existing language annotations with non-existent factors while controlling other variables. For example, “go push the blue block” becomes “go push the orange block” for a visual defect. Similarly, we modify adjectives, nouns, and verbs to create defects in other dimensions. Subsequently, the dataset size is expanded significantly by incorporating synonymous expressions. Finally, we add 7.3k instructions containing multiple variable perturbations, where multiple variables are replaced within a single instruction. These two steps aim to enhance the dataset’s diversity and generalizability. For safety and out-of-context defective instructions, we meticulously craft prompts for GPT-4o [6] to autonomously generate these instructions.
We then convert the language annotations into the chat-like expression of MLLM assistants. CALVIN contains one language instruction and a list of (observation, action) pairs $[x_{task},(o_{0},a_{0},o_{1},a_{1},...,o_{t},a_{t},...)]$ per trajectory. We programmatically generate texts in the format of chat interactions using templates. A simple example of user-assistant interactions, is “User: Can you help me $x_{task}$? Assistant: Sure, I will $x_{task}$ <ACT>." For defective instructions, it will be “User: Can you help me $x_{task}$? Assistant: Sorry, I can not $x_{task}$ <REJ>." This process trains the model to recognize and appropriately respond to everyday human instructions, making informed decisions on whether to execute or refrain from executing actions. It thereby fosters a conversational interface capable of seamlessly transitioning from dialogue to actionable responses.
III-D Evaluation
We have developed a fair evaluation protocol and success criterion for our benchmark. The rollout begins with a defective instruction, followed by the CALVIN LH-MTLC tasks [17] with complex, unseen annotations, designed to assess the impact of defective instructions over the entire sequence and language generalization. To ensure fairness, we exclude defective tasks from success rate calculations and design them not to affect the proper initiation of subsequent tasks. All other settings and metrics adhere to the CALVIN evaluation framework.
Figure 4: The comparison between different methods to tackle RAMA. Top: Directly fine-tune a VLA model to handle various instructions. Middle: Use a pre-trained VLM as the high-level identifier to judge various instructions. If the task is executable, then transmit it to the low-level policy. Bottom: The dual-system VLA combines high-level VLM and low-level policies through the action latent embedding. In this way, the model inherits the language understanding capabilities of the pretrained VLM and transmits the latent action representation to the downstream policy without loss, enabling it to produce correct actions under various instructions.
Figure 5: Overview of RationalVLA. Given the observation and text instruction including the defective one, an MLLM generates a text description of a task and an <ACT> token or a <REJ> token. The latent space embeddings from the <ACT> token’s last layer serve as a high-level latent goal for the downstream policy and those of <REJ> token reject defective instructions in the current scene. Our dual-system model can run the high-level MLLM reasoning and low-level action policy execution loops asynchronously. During inference, the action policy frequently updates actions based on environment changes and the latest <ACT> token’s embedding, while the MLLM updates less frequently, enabling efficient inference.
IV RationalVLA
As previously discussed, existing models face challenges in effectively handling defective instructions. To address this, we propose a novel model RationalVLA.
IV-A Comparison of Architectures
Fine-tuned VLA. Given the strong generalization capabilities on visual and language modalities demonstrated by pretrained VLAs, a natural approach is to fine-tune an end-to-end VLA model to process and understand complex instructions. However, several key challenges limit its performance: (1) Fine-tuning a VLA model often leads to catastrophic forgetting [18], resulting in degraded language capabilities, e.g., RoboFlamingo reaches a mediocre performance on the enriched instructions. (2) Considering downstream robotic tasks, VLA models frequently underperform compared to specialized expert policies, e.g., RoboFlamingo lags behind 3DDA on standard CALVIN benchmarks.
VLM as Identifier. To address the aforementioned issues, a straightforward idea is to leverage both the multimodal reasoning capabilities of pretrained VLMs and the manipulation proficiency of a task-specific policy. Specifically, a separately trained VLM can serve as a high-level identifier that decides whether to reject a given instruction, after which a downstream policy is responsible for task execution. This constitutes a rudimentary dual-system architecture. With the capabilities of powerful VLMs such as GPT-4o, this design can effectively filter out inappropriate instructions. However, such a naïve combination fails to fully exploit the VLM’s language understanding capabilities. When faced with complex instructions, the downstream policy often lacks sufficient comprehension, leading to task failure. Consequently, this approach falls short of achieving our overarching goal: aligning the robotic agent’s action space with the complexity of natural language instructions.
Dual-System VLA. To fully leverage the language understanding capabilities of pretrained VLMs and the manipulation proficiency of pretrained policies, we propose a Dual-System VLA architecture, which seamlessly bridges the gap between the language space and the action space. In this framework, the action representations produced by the VLM are mapped to a shared latent embedding space, which subsequently serves as a conditioning signal for the low-level policy. During end-to-end fine-tuning, the latent embedding remains learnable, ensuring alignment between the high-level output space and the low-level input space. This process enables the accurate transmission of holistic action goals and subtle intent to the low-level policy, as these aspects are often difficult for the low-level policy to understand through language alone.
Specifically, we implement this latent transmission by introducing only two special tokens into the vocabulary. This minimal design choice preserves the original embedding space for language tokens and helps prevent catastrophic forgetting during joint training. As a result, our method can effectively handle complex and potentially ambiguous instructions, thereby supporting robust task execution across diverse scenarios. We provide a detailed description of our dual-system model in the following section.
IV-B Our Dual-system VLA
RationalVLA unifies the capabilities of a slow but powerful pre-trained MLLM with a fast and simple decision-making policy to create a model that ingests vision, language, and proprioception inputs to output low-level actions. This integration involves a dual system: a pre-trained MLLM $f_{\phi}$ and a pre-trained policy $\pi_{\theta}$, which are parameterized by $\phi$ and $\theta$ respectively.
High-level MLLM. The MLLM $f_{\phi}$ comprises a large language model $f_{\textnormal{LLM}}$ and a vision encoder $f_{\textnormal{encoder}}$, which can project images into the text-only large language model embedding space, thereby facilitating a multimodal understanding of textual and visual inputs.
In our framework, given the third-view RGB image $I$, the vision encoder $f_{\textnormal{encoder}}$ first encodes it into image features, and then the projector $\varPhi$ maps the features into the visual tokens $h_{I}$ in the LLM input space. Along with the input image, the text instructions $S$ are tokenized into text tokens $h_{S}$ by the LLM tokenizer $T$. Then the $f_{\textnormal{LLM}}$ takes in text tokens $h_{S}$ and image tokens $h_{I}$ and autoregressively generates text response $y_{S}$. The whole process can be formulated as:
$y_{S}$ $=f_{\phi}(S,I)=f_{\textnormal{LLM}}(h_{S},h_{I})$ (1)
$=f_{\textnormal{LLM}}(\varPhi(f_{\textnormal{encoder}}(I)),T(S)).$
We expand the vocabulary of the language model with an additional <ACT> token and a <REJ> token. The <ACT> token extracts the vision-language information, representing the latent space embedding of the action entity. Upon generating <ACT> token, its hidden embedding is further fed into the low-level policy as a condition. Some previous works have focused on dynamically processing neural network inputs [19, 20], which is equally effective for robots to flexibly handle language instructions. The model is trained to output either <ACT> tokens $y_{<ACT>}$ or <REJ> tokens $y_{<REJ>}$ depending on whether the instructions are executable or not in the current scene. Then, the $y_{<ACT>}$ is projected by a projector $\varPsi$ and further unsqueezed in the zero dimension to align with the policy latent conditioning space. Thus, we get the latent embeddings $z=unsqueenze(\varPsi(y_{<ACT>}))$. For $y_{<REJ>}$, it is projected to be $z_{0}=\varPsi(y_{<REJ>})$, which is an all-zero tensor to stop action. Finally, $z$ is fed into the policy $\pi_{\theta}$.
Low-level Policy. We employ 3D Diffuser Actor [16] as low-level Policy $\pi_{\theta}$. It takes as input user instructions $S$, visual observations at the current timestep $o$, and proprioception $p$, and output actions at the current timestep $a$:
$a=\pi_{\theta}(o,p,z).$ (2)
Here we modify it to receive the conditioning latent embeddings $z$ in replace of $S$ in order to better bridge the MLLM’s response and the low-level policy. Each action $a$ describes an end-effector pose and is decomposed into 3D position, 3D orientation and a binary open/closed state: $a=\{a^{\mathrm{\small pos}}\in\mathbb{R}^{3},a^{\mathrm{\small rot}}\in\mathbb{R}^{6},a^{\mathrm{\small open}}\in\{0,1\}\}$. We train a denoising neural network that takes as input a 3D scene visual encoding, the current estimate of the end-effector’s future trajectory, as well as the diffusion iteration index, and predicts the error in 3D translation and rotation. It marries diffusion policies for handling action multimodality and 3D scene encodings for effective spatial reasoning.
[…]
Architecture Method Test set Success Rate (%) Average
1/5 2/5 3/5 4/5 5/5 length
Single Policy 3DDA[16] D (unseen, defective) 43.7 21.5 12.2 7.7 3.0 0.88
Single VLA RoboFlamingo[3] D (unseen, defective) 46.0 28.3 21.0 10.7 5.0 1.18
VLM as Identifer LLaVA [5] + 3DDA D (unseen, defective) 59.8 32.2 15.7 8.3 4.9 1.21
VLM as Identifer GPT-4o [6] + 3DDA D (unseen, defective) 64.3 33.1 16.0 10.0 6.2 1.30
Dual-system VLA RationalVLA (ours) D (unseen, defective) 74.3 58.3 42.3 30.0 20.7 2.26
The prediction of <REJ> token can be seen as a VQA task, where the spatial relationship in the scene and the feasibility of the instructions should be considered. Thus, training <REJ> token further improves the quality of multimodal reasoning and ameliorates the hallucination of MLLM.
IV-C Training
The training of RationalVLA employs a combination of techniques to integrate the MLLM and policy components. We leverage Low Rank Adaptation [52] (LoRA) for the supervised fine-tuning of the MLLM, allowing for more efficient training.
Stage 1: Pre-training for feature alignment. We adopt a cold start approach to policy training, reminiscent of staged training strategies seen in prior works, by first freezing the low-level policy and only fine-tuning the language model. This preliminary phase focuses on aligning the embeddings produced by the MLLM with the feature space of the policy, to prevent the initial unstable gradient due to the mismatch between the pre-trained policy and the pre-trained LLM.
Stage 2: Fine-tuning end-to-end. In this stage, we unfreeze the low-level policy and proceed with fine-tuning the entire model in an end-to-end manner. This phase constitutes the majority of the training process.
Loss Function. The loss function is comprised of 3 terms, and can be expressed as follows:
$\mathcal{L}=$ $\lambda_{1}\mathcal{L}_{\textnormal{policy}}(\pi_{\theta},o_{t},z_{{\texttt{<ACT>}}},a_{t})+\lambda_{2}\mathcal{L}_{\textnormal{LLM}}(f_{\phi},x_{\textnormal{txt}},x_{\textnormal{img}})$ (3)
$+\lambda_{3}\mathcal{L}_{\textnormal{CLIP}}(z_{{\texttt{<ACT>}}},g_{\textnormal{txt}}),$
where $\mathcal{L}_{\textnormal{policy}}(\cdot)$ is the diffusion denoise loss of low-level policy [16]. $\mathcal{L}_{\textnormal{LLM}}(\cdot)$ is the same auto-regressive training objective following [5], with a special token <ACT> instead. $\mathcal{L}_{\textnormal{CLIP}}(z_{\textnormal{act}},g_{\textnormal{txt}})$ is used to regularize the latent embedding $z_{{\texttt{<ACT>}}}$, ensuring it well aligned with the lower level ground truth text description $g_{txt}$ [39]. Specifically, we employ $\mathcal{L}_{\textnormal{CLIP}}(z_{\textnormal{act}},g_{\textnormal{txt}})=1-\cos(\textnormal{stop\_gradient}(\textnormal{clip}(g_{\textnormal{txt}})),z_{\textnormal{act}})$. This auxiliary loss helps regularize the predicted embedding. We set $\lambda_{1}=1$, $\lambda_{2}=100$, and $\lambda_{3}=1$. The $\lambda_{2}$ owns a high value because the LLM necessitates extensive finetuning due to the newly designed special tokens.
V Experiments
In this section, we conduct comprehensive experiments to validate the efficacy of our method.
Implementation Details. During training, we use LoRA with a rank of 16. All the reported results employ the pre-trained LLaVA-v1.5-7b [5] as VLM $f_{\phi}$ and CLIP-VIT-L/14-336 [53] as vision encoder $f_{enc}$. The implementation of low-level policy follows [16]. The fine-tuning data is a mixture of the CALVIN A, B, and C splits and the RAMA training set, combined in a ratio of 0.7 to 0.3, respectively. The fine-tuning of RationalVLA involves two stages: 5,000 iterations in the first stage and 30,000 in the second. The total training process takes 36 hours on eight 80GB H800 GPUs.
Metrics. Following CALVIN [17], we report success rates (%) along with the average length of completed tasks (out of the whole 5 tasks) per evaluation sequence.
V-A Baselines
Fine-tuned policy. 3D Diffuser Actor (3DDA) [16] serves as our low-level policy. For comparison, we also fine-tune it on the RAMA dataset to produce static actions in response to defective instructions, thereby preventing any negative impact on subsequent tasks.
Fine-tuned VLA model. With inherited language abilities, we fine-tune a typical VLA model, RoboFlamingo, to output no in response to defective instructions.
VLM as identifier. We seperately employ the popular open-sourced VLM LLaVA [5] and the most advanced proprietary model GPT-4o to form LLaVA + 3DDA and GPT-4o + 3DDA. Notably, since the combination of two separate models cannot facilitate gradient backpropagation, they are not fine-tuned on the RAMA dataset. Instead, we make a prompt engineering to leverage their inherent multimodal perception capabilities. The prompt is shown in Fig.4.
V-B Rational Manipulation
Settings. To evaluate the capabilities of handling diverse instructions in real-life scenarios, we test models on the unseen CALVIN D split with unseen instructions [3] in the RAMA evaluation protocol, following subsection III-D. This setting poses challenges to the language generalization and multimodal reasoning capabilities of models.
Quantitative Results. The quantitative results are shown in Table II. The 3DDA achieves a low success rate primarily because its language encoder provides a limited language ability, rendering it incapable of comprehending unseen instructions or distinguishing defective instructions. The LLaVA + 3DDA system achieves exceptionally high first-task success rates. This stems from its well-designed prompts that guide the pretrained VLM to first assess the executability of language instructions. This design fully unleashes the low-level action policy’s manipulation capabilities, ensuring it operates without interference from defective instructions. However, some unseen instructions and defective instructions are difficult to distinguish, requiring in-depth understanding and reasoning. LLaVA occasionally confuses these instructions. GPT-4o shows substantial improvements in both multimodal perception and language reasoning capabilities, resulting in reduced susceptibility to defective instructions and higher first-task success rates. Nevertheless, for the VLM-as-identifier system, while the high-level VLM can help the low-level policy reject defective instructions, it does not directly enhance the language generalization capabilities of the system. This is because the VLM can not directly translate unseen complex instructions for policies and further execute corresponding actions. Consequently, its advantages in subsequent task success rates and average trajectory length remain marginal.
Compared with the best baseline GPT4o + 3DDA, our method RationalVLA yields an improvement from 1.30 to 2.26 in average length. The success rate of the last task is elevated by 14.5%. RationalVLA learns a <REJ> token to enable rejections for non-executable tasks, which enables rejections for non-executable tasks. Thus, the appearance is little affected by defective instructions. In the following tasks, the high-level VLM translates complex, unseen instructions into an understandable latent action embedding, which is then taken as input by the low-level policy to generate effective actions.
Figure 6: Experimental results of our RationalVLA on the different dimensions of RAMA benchmark.
Different Dimensions. Fig. 6 illustrates the performance of our RationalVLA across different dimensions of RAMA benchmark. Our RationalVLA achieves particularly strong results on the out-of-context dimension, which can be attributed to the VLM’s robust language capabilities. In terms of visual perturbations, the VLM’s spatial perception capability is further enhanced through training on the RAMA dataset, reducing the visual hallucinations. However, instructions with motion perturbation are often highly deceptive, resulting in the lowest average trajectory length.
Figure 7: Visualization of the rejection mechanism. The defective instructions are colored in burgundy. At the right of every task, we denote the success or failure. The first row shows a failure case of the fine-tuned RoboFlamingo, it fails and terminates in task 1, with a total execution time of 76s. This perturbation makes the model inefficient and incapable of completing sequential tasks. The rest of the rows show a successful case of our RationalVLA , it successfully finishes all the tasks and uses only 48s, showing the efficiency when tackling multiple tasks.
Visualization and Qualitative Analysis. Fig. 7 displays a failed rollout of the fine-tuned RoboFlamingo and a successful rollout of our RationalVLA . When receiving defective instructions, RoboFlamingo assumes the red block instead of the orange block as the target, then lifts it and puts it in the cabinet. This makes the robot unable to lift the red block from the table in the next task, thus failing in task 1. The whole process takes 76s because the model must expire the limited timesteps to find itself lost. The existence of defective instructions affects the normal action output and disrupts the layout of the environment, leading to inefficient processes. Instead, when receiving the instruction, our VLM part successfully recognizes that the orange block was absent in the environment and immediately outputs a <REJ> token. This <REJ> token halts the current task, allowing the system to transition seamlessly to the next task. Then RationalVLA completes the whole rollout successfully. The entire process is highly efficient, with the <REJ> token playing a crucial role. This also demonstrates the performant manipulation capability of our RationalVLA.
V-C Manipulation with Unseen Instructions
Settings. To evaluate the capabilities of handling diverse instructions in real-life scenarios, we test models on the CALVIN D split with unseen instructions [3] in the RAMA evaluation protocol.
Results. While RoboFlamingo leverages pretrained vision-language alignment, it struggles with nuanced instruction understanding and accurate task grounding under this enriched setting III. Similarly, VLM-as-Identifier architectures fail to distinguish unseen instructions, often misclassifying them or defaulting to irrelevant behaviors.
In contrast, our method, RationalVLA, combines the MLLM for instruction interpretation with a robust multitask policy. This design enables the model to semantically parse unseen instructions and accurately associate them with corresponding actions. Notably, RationalVLA achieves an average gain of 1.05 average length over the strongest baseline, significantly outperforming all comparison methods in this setting. These results demonstrate superior generalization and interpretability in enriched manipulation tasks.
[…]
Architecture Method Test set Task completed in a row (%) Average Length
1 2 3 4 5
Single VLA RoboFlamingo D 82.4 61.9 46.6 33.1 23.5 2.48
Single Policy 3DDA [16] D 93.8 80.3 66.2 53.3 41.2 3.35
VLM as Identifier GPT-4o [6] + 3DDA [16] D 93.8 80.3 66.2 53.3 41.2 3.35
Dual-system VLA RationalVLA D 95.7 83.0 68.3 58.0 48.0 3.53
Single VLA RoboFlamingo D (unseen) 62.0 33.0 16.4 8.6 4.6 1.25
Single Policy 3DDA [16] D (unseen) 65.2 39.1 20.3 11.7 6.1 1.42
VLM as Idetifier GPT-4o [6] + 3DDA [16] D (unseen) 65.2 39.1 20.3 11.7 6.1 1.42
Dual-system VLA RationalVLA D (unseen) 80.9 63.1 46.7 34.1 23.6 2.48
V-D Manipulation With Seen Instructions
Settings. To demonstrate the effectiveness of our method on conventional manipulation tasks, we evaluate it on the CALVIN ABC→D split with seen instructions only. This setting highlights that our model outperforms baselines not only in challenging scenarios but also under traditional task distributions.
Results. As shown in Table III, RationalVLA achieves the highest 48% 5/5 task success rate and an average correct step length 0.18 higher than the strongest baseline [16]. This result demonstrates that RationalVLA remains highly competitive on standard manipulation benchmarks. Importantly, it not only retains the control effectiveness of the underlying policy, but also enhances task performance by leveraging the VLM’s language understanding and multimodal reasoning through seamless integration.
V-E Ablation Study
Both settings of the ablation study follow subsection V-B.
Clip Loss during Training Process. Table IV shows the performance of RationalVLA drops 1.22 in average length without clip loss. This emphasizes the necessity of clip loss for aligning the latent space embedding and the language input of the pre-trained low-level policy during supervised fine-tuning.
Chat-like expression. Table IV shows that the performance of RationalVLA improves 0.78 in average length through chat-like expression. This indicates that chat-like expressions are closer to the original language distribution, preventing catastrophic forgetting [54] and maintaining the inherent performance of the language model.
TABLE IV: Ablation results on important designs of RationalVLA.
Task completed in a row (%) Average 1 2 3 4 5 Length w/o clip loss 51.0 27.3 13.3 7.7 4.7 1.04 w/o expression 59.3 41.0 26.3 13.0 8.3 1.48 RationalVLA 74.3 58.3 42.3 30.0 20.7 2.26
VI Real-world Experiments
To further validate the practical performance and generalization on language-conditioned tasks, we validate our model in the real world.
VI-A Robotic System Setup
We built a teleoperated robot system to collect high-quality demonstrations and conduct experiments. The system includes a lead arm and a follower arm. The follower arm consists of a 6-DOF AgileX Piper robotic arm with a 1-DOF gripper and an ORBBEC Dabai depth camera for first-view image. We also utilize a RealSense D515 camera to capture third-view images. The other Piper arm with a teaching gripper serves as the lead arm. During data collection, the lead arm is controlled by a human expert, and the follower arm mirrors the same action. For each task with different levels of difficulty, we collect over 25 demonstrations through human operation, accompanied by annotated instructions. During the test, only the follower arm is activated, and each task is tested over 20 episodes.
Baselines. We choose our low-level policy 3DDA and the strongest combination GPT-4o + 3DDA as our baselines.
VI-B Basic Tasks With Unseen Instructions
We meticulously designed three basic tasks with unseen instructions. These tasks include “Lift the watermelon”, “Put the pepper in the bowl”, and “Stack green bowl on yellow bowl”, but their instructions are replaced with complex, unseen ones, e.g., “Elevate the Citrullus lanatus specimen”.
Results. We report success rates (%) in Table V and trajectory visualizations in Fig 9, which demonstrate that RationalVLA generalizes effectively to real-world manipulation tasks with unseen instructions. Despite the increased complexity and variability of real-world scenes, RationalVLA maintains strong performance, achieving a notable 68.3% improvement in success rate over 3DDA. This highlights the strength of integrating an MLLM capable of interpreting enriched instructions with a robust control policy, preserving both visual understanding and language generalization in practical deployment.
Figure 8: Real-world robotic system setup. We employ a 6-DOF AgileX Piper arm equipped with a 1-DOF gripper and utilize a RealSense D515 camera to capture third-view images.
TABLE V: Success rates (%) on tasks and average success length with unseen instructions in the real world.
Model Lift Put Stack Average
3DDA [16] 25 20 5 16.7
GPT-4o [6] + 3DDA [16] 25 20 5 16.7
RationalVLA (ours) 90 75 90 85.0
TABLE VI: Success rates (%) on twofold tasks and average success length with defective instructions in the real world.
Model Lift-put Put-put Put-stack Average 3DDA [16] 0 0 0 0.0 GPT-4o [6] + 3DDA [16] 15 5 5 8.3 RationalVLA (ours) 80 70 90 80.0
Figure 9: Representative results on defective instruction and unseen instruction of real-world experiments. The sequential images showcase that the robotic arm rejects the defective instructions ,understands unseen instructions and completes the next task.
VI-C Long-Horizon Tasks
We designed three long-horizon tasks to evaluate the impact of defective instructions in a long sequence. Each task comprises two subtasks: the first task with defective instructions and the second task with unseen instructions.
  • Lift-put: “Lift the watermelon” serves as the defective instructions with no watermelon on the scene. Then the model should finish “put the pepper into the bowl”.
  • Put-put: The defective instruction is “Put the grape in the cup”.However, there is no cup on the scene. The second task is “put the watermelon into the bowl”.
  • Put-stack: The defective instruction is “Put the watermelon into the yellow bowl”, but the watermelon lies outside the robot’s reachable workspace, making the task infeasible. The second instruction is “Stack the yellow bowl on the green bowl”.
Results. We report success rates in Table VI and trajectory visualizations in Fig 9, which demonstrate that our RationalVLA effectively generalizes to real-world manipulation tasks even under defective instructions. Unlike 3DDA, which relies solely on a policy-based architecture and often executes all instructions indiscriminately, RationalVLA learns to reject infeasible commands in the current environment, thereby preserving the integrity of subsequent tasks. This results in an average 63.7% improvement in success rate, underscoring the benefits of incorporating instruction comprehension into the control pipeline for robust real-world deployment.
VII Conclusion
This paper analyzes the limitation of the classic manipulation task and proposes a new benchmark, RAMA, which allows defective instructions. It contains more than 14k data samples. To better manage RAMA benchmark, we propose RationalVLA. It is a dual-system vision-language-action model with the capability of processing complex instructions with defective ones. Extensive experiments show that our model demonstrates surprisingly effective performance in both classic manipulation tasks and rational manipulation tasks. RAMA benchmark research further highlights its value in industrial and domestic applications. We hope our work will accelerate the deployment of robot assistants in real-world applications.
Appendix A Acknowledgment
We sincerely thank Yuxin Huang for creating the exquisite video. Pengxu Hou assisted with hardware setup. We are grateful to Chris McCool for his valuable discussions and suggestions.
Figure 10: Defective Instruction Generator for RationalVLA. Our defective instruction generator is divided into two parts, Modular Generator and Direct Generator. The Modular Generator is designed to produce defective instructions for the visual, physical, semantic, and motion dimensions by systematically replacing specific linguistic components (Adj., Noun., Verb.) while maintaining strict control over variables. The Direct Generator focuses on generating defective instructions of the safety and out-of-context dimensions, leveraging carefully crafted prompts to query GPT-4o.
Appendix B Data
B-A Instruction Generator
The defective instruction generator used for RationalVLA  is shown in Figure 10. Our defective instruction generator is divided into two parts, Modular Generator and Direct Generator.
Modular Generator is used for generating the visual, physical, semantic, and motion defective instructions. We begin by analyzing the existing instructions from the CALVIN environments to identify different linguistic parts. This stage is completed by GPT-4o [6] and the prompt used here is shown in Figure 12. Subsequently, we programmatically replace the existing language annotations from CALVIN with non-existing factors. Other variables remain the same. These factors are randomly selected from different libraries and replace the corresponding linguistic components (Adj., Noun., Verb.) of annotations. Specifically, for the visual, physical, semantic, and motion dimensions, defective instructions are generated by systematically replacing specific linguistic components in the original annotations. Adjectives are replaced to introduce defects in the visual and physical dimensions, nouns are substituted to address the semantic dimension, and verbs are modified to create defects in the motion dimension. This approach ensures a controlled and consistent methodology for generating defective instructions across all dimensions.
Direct Generator is used for generating the safety and out-of-context defective instructions, as shown in Figure 10. This generator leverages meticulously designed prompts to query GPT-4o, enabling the creation of instructions tailored for these dimensions. Our approach successfully produces defective instructions that meet the desired requirements, largely because these two dimensions do not necessitate the stringent variable control required by the other four dimensions.
In our experiments, the instructions generated by GPT-4o are unable to fully satisfy the specific control variable requirements for the visual, physical, semantic, and motion dimensions. Therefore, we opt to design tailored codes to handle this process, ensuring precise control and consistency across these dimensions. This division ensures both precision and adaptability in creating diverse and representative defective instructions across all dimensions.
B-B Data Examples
The RAMA dataset encompasses defective instructions across six dimensions: visual, physical, semantic, motion, safety, and out-of-context. Examples illustrating each of these dimensions are provided in Figure 11. To enhance the diversity and generalizability of RAMA benchmark, we extend the original dataset with mixed instructions that involve multiple variable perturbations.
Figure 11: Some examples of defective instructions in RAMA benchmark for Defective Instruction Generator.
B-C Prompt
The GPT-4o prompt used for analyzing part of speech in instructions is shown in Figure 12. The results from this analysis are subsequently leveraged to modify variables for generating defective instructions across each dimension. The GPT-4o prompts used for autonomously generating safety & out-of-context instructions are shown in Figure 13. Prompts here are inspired by [55].
Figure 12: Prompt for analyzing the linguistic parts of instructions.
Figure 13: Prompt for generating defective instructions of safety & out-of-context dimensions.
B-D Chat Templates
The chat templates used for RationalVLA are shown in Figure 14. The system prompt follows the type of llama-v0 [56]. We randomly replace instructions in CALVIN with defective instructions in RAMA benchmark. We design templates for questions and answers as well.
Figure 14: Chat templates for RationalVLA training process.